The following events apply to Lewis and Harper, a public accounting firm, for the Year 1 accounting period:

  1. Performed $82,000 of services for clients on account.
  2. Performed $48,000 of services for cash.
  3. Incurred $44,000 of other operating expenses on account.
  4. Paid $19,000 cash to an employee for salary.
  5. Collected $62,000 cash from accounts receivable.
  6. Paid $36,000 cash on accounts payable.
  7. Paid a $6,000 cash dividend to the stockholders.
  8. Accrued salaries were $3,600 at the end of Year 1.


Required
a. Show the effects of the events on the financial statements using a horizontal statements model like the following one. In the Cash Flow column, use OA to designate operating activity, IA for investment activity, FA for financing activity, NC for net change in cash and NA to indicate the element is not affected by the event. The first event is shown as an example. (Enter any decreases to account balances and cash outflows with a minus sign.)
b. What is the amount of total assets at the end of Year 1?
c. What is the balance of accounts receivable at the end of Year 1?
d. What is the balance of accounts payable at the end of Year 1?
f. What is net income for Year 1?
g. What is the amount of net cash flow from operating activities for Year 1?

Show the effects of the events on the financial statements using a horizontal statements model like the following one. In the Cash Flow column, use OA to des investment activity, FA for financing activity, NC for net change in cash and NA to indicate the element is not affected by the event. The first event is shown as decreases to account balances and cash outflows with a minus sign.) LEWIS AND HARPER Statements Model For Accounting Year 1 Stockholders Equity Retained Assets Liabilities Income Statement Event No. Statement of Cash Flow Accounts Receivable Accounts able Salaries + able Cash+ Revenue-| Expense Net Income Earnin 82,000 = 82,000 82,000- 82,000 NA 2 4 Totals Req A Req B to FReq A Req B to F Req G What is the amount of net cash flow from operating activities for Year 1? Net cash flow from operating activities for Year 1 < Req B to F Req G
Answer #1

The answer has been presented in the supporting sheets. All the parts has been solved with detailed explanation. For complete solution refer the supporting sheets.

ssssssss - Microsoft Excel Home Insert Page Layout Formulas Data Review View Cut Σ AutoSum Aa ,I 11 wrap Text (General Copy Paste r Format Painter B 1 u. 田 Cr -- = 函Merge & Center- al;o Corr ditti nalasoamat . Styles. s % , Insert Delete Format Sort &Find & - 2 ClearFilter Select Editing Formatting as Table Styles Clipboard Font Alignment LEWIS AND HARPER STATEMENTS MODEL For Accounting Year 1 Amount in S)- Stockholders Assets Liabilities Income Statement Statement of Cash Salaries Retained Cash Revenue | - | Expense recievablePayable Payable earnings NA 48000 OA NA 19000 OA 62000 OA 36000 OA 6000 FA NA 49000 NC 820001 82000 8 82000 820001 48000+ 44000 440001=1-44000 19000 19000+ 62000 + 36000 -6000+ 19000 19000 -620001 = 12 36000 + 13 3600t 3600 36001 = 3600 15 Totals 16 17 18 19 49000+ 20000 - 8000 + 3600+ 57400 130000 66600 63400 21 23 l4 ‘ト커 Sheet1 Sheet2 .. Average: 4966666667 Count: 17 Sum: 447000 | MU-105% 10:21 AM Search the web and Windows

(S - 469 Bs Ranicahon 2 Paid ecttusbles 60o 55000